# Graph Attention-Driven Bayesian Deep Unrolling for Dual-Peak Single-Photon Lidar Imaging

### v1

@@ V-A Results on synthetic data
Quantitative comparison. Table I and II report the quantitative results for the Art scene and the Bowling scene, respectively. To analyze the results, we report the Chamfer distance with two terms separately: O-To-G (distance from the estimated point cloud to the ground truth) and G-To-O (distance from the ground truth to the estimated point cloud). In all scenarios, the proposed method shows the lowest error for G-To-O, which indicates that the estimated point cloud is closer to the ground truth, albeit with some outliers. When PPP and SBR are both 1, Peng’s method shows the lowest errors except for G-To-O, but as PPP and SBR increase, the proposed method shows the better performance overall. Fig. 7 visualizes the errors with different levels of PPP and SBR for the Art scene. When PPP and SBR are both 1, Peng’s method shows the lowest error. As PPP and SBR increase, the proposed method shows the lowest error overall. Table III compares the running time of the proposed method with other methods. While other methods run twice, the proposed method only once. The runtime (13.5 seconds) of the proposed method consists of two parts: estimating the initial multiscale (11.49 seconds) and the inference time (2.8 seconds). The number of parameters of the proposed method is shown to be the smallest among the methods, which indicates the efficiency of the proposed method.
Ablation study. To analyze the network architecture, we perform the ablation study in terms of the number $L$ of multiple scales and the number of graph attention layers. As shown in Table IV, the number of multiple scales affects the performance. We choose $L=4$, because increasing $L$ beyond this value does not significantly improve the results. We also report the effect of the number of graph attention layers. This number includes all the graph attention layers in the squeeze part, the expansion part for each stage. Since the number of graph attention layers affects the number of parameters and the computational cost, we choose 45 graph attention layers, as a trade-off between performance and computational cost.
@@
L Convolutions k DAE Chamfer dist. Parameters
2 45 6 0.0595 0.0293 17808
3 0.0069 0.0040 19416
4 0.0067 0.0040 21024
5 0.0072 0.0037 22632
4 15 0.0069 0.0040 1056
30 0.0069 0.0040 7584
45 0.0067 0.0040 21024
SBR$=8.57$(original)
@@
Peng (2 runs) RT3D BU3D (2 runs) Proposed Uncertainty
@@ V-B Results on real data
We evaluate the proposed method on the real dual-peak data of Mannequin behind scattering object [46] with different SBR levels. The histogram dataset has a spatial-time resolution of $99\times 99\times 4001$ with PPP=45 and SBR=8.57. To generate different SBR levels, we add background photons, so that the new SBR levels are 8.57, 2.14, 1.07, 0.54, and 0.27. For the comparison, we run Peng [27] and BU3D [31] two times in the region of scattering medium and the Mannequin, separately. In this regard, they have the advantage of knowing the approximate positions of the surfaces. We additionally compare our results to a multi-surface reconstruction method called RT3D [18]. As RT3D supports multiple surfaces per pixel, it is run only once.
The reconstruction results are shown in Fig. 8. As the SBR level decreases, the quality of the reconstructed point clouds decreases for other methods. When SBR=0.27, the RT3D fails to reconstruct meaningful point clouds. The proposed method outperforms the other methods in terms of the quality of the reconstruction. As the SBR decreases, we observe the higher uncertainty values. This experiment shows the effectiveness of the proposed method in the real dataset with an accurate reconstruction having uncertainty estimation.
@@ VI Conclusion
We have proposed a novel deep learning method for dual-peak reconstruction in single-photon Lidar imaging. Unlike existing deep learning-based methods, the proposed method supports the dual-peak reconstruction in a single run with uncertainty information. The experimental results show that the proposed method presents less artifacts such as bleeding effects with accurate results. A limitation of the proposed method is that it often observes outliers in low photon cases. As a future direction, we consider extending from dual-peak to multi-peak reconstruction and incorporating reflectivity information in the model.

### v2

@@ V-A Results on synthetic data
Quantitative comparison. Table I and II report the quantitative results for the Art scene and the Bowling scene, respectively. To analyze the results, we report the Chamfer distance with two terms separately: O-To-G (distance from the estimated point cloud to the ground truth) and G-To-O (distance from the ground truth to the estimated point cloud). In all scenarios, the proposed method shows the lowest error for G-To-O, which indicates that the estimated point cloud is closer to the ground truth, albeit with some outliers. When PPP and SBR are both 1, Peng’s method shows the lowest errors except for G-To-O, but as PPP and SBR increase, the proposed method shows better performance overall. Fig. 7 visualizes the errors with different levels of PPP and SBR for the Art scene. When PPP and SBR are both 1, Peng’s method shows the lowest error. As PPP and SBR increase, the proposed method shows the lowest error overall. Table III compares the running time of the proposed method with other methods. While other methods run twice, the proposed method only once. The runtime (13.5 seconds) of the proposed method consists of two parts: estimating the initial multiscale (11.49 seconds) and the inference time (2.8 seconds). The number of parameters of the proposed method is shown to be the smallest among the methods, which indicates the efficiency of the proposed method. The proposed method requires 20.3 GB of GPU memory for training. For testing on the Art scene, it requires 21.5 GB for initial mulstiscale estimation and 1.7 GB for inference.
Ablation study. To analyze the network architecture, we perform an ablation study in terms of the number $L$ of multiple scales and the number of graph attention layers. As shown in Table IV, the number of multiple scales affects the performance. We choose $L=4$, because increasing $L$ beyond this value does not significantly improve the results. We also report the effect of the number of graph attention layers. This number includes all the graph attention layers both in the squeeze parts and the expansion parts. Since the number of graph attention layers affects the number of parameters and the computational cost, we choose 36 graph attention layers, as a trade-off between performance and computational cost. The reported uncertainty corresponds to the average uncertainty of the depth absolute error (DAE) across all pixels. We also studied an effect on two hyperparameters: the number of neighbors in the k-NN and the Gumbel-Softmax temperature used for hard attention, where we observed no significant performance changes. Throughout the experiments, we set k in kNN to 6 and the temperature to 10.
@@
L #Conv. k in kNN Temperature DAE Uncertainty #Params
2 36 6 10 0.0595 0.0258 17808
3 0.0069 0.0319 19416
4 0.0067 0.0078 21024
5 0.0072 0.0151 22632
4 12 6 10 0.0069 0.0104 1056
24 0.0069 0.0087 7584
36 0.0067 0.0078 21024
4 36 4 10 0.0066 0.0078 21024
6, 8 0.0067 0.0078 21024
4 36 6 0.1-100 0.0067 0.0078 21024
SBR=8.57=8.57
@@
RT3D Peng (2 runs) BU3D (2 runs) Proposed Uncertainty
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
RT3D [18] Lindell [25] (2 runs) Peng [27] (2 runs) BU3D [31] (2 runs) Proposed Uncertainty
Figure 9: Reconstructed point clouds from real dataset [25] modified to contain dual peaks. Lindell, Peng and BU3D are run twice to estimate the first and second surfaces separately. RT3D and the proposed method.
@@ V-B Results on real data
We evaluate the proposed method on the real dual-peak data of Mannequin behind scattering object [46] with different SBR levels. The histogram dataset has a spatial-time resolution of $99\times 99\times 4001$ with PPP=45 and SBR=8.57. To generate different SBR levels, we add background photons, so that the new SBR levels are 8.57, 2.14, 1.07, 0.54, and 0.27. For the comparison, we run Peng [27] and BU3D [31] two times in the region of scattering medium and the Mannequin, separately. In this regard, they have the advantage of knowing the approximate positions of the surfaces. We also compare our results to the multi-surface reconstruction method RT3D [18].
The reconstruction results are shown in Fig. 8. As the SBR level decreases, the quality of the reconstructed point clouds decreases for other methods. When SBR=0.27, the RT3D fails to reconstruct meaningful point clouds. The proposed method outperforms the other methods in terms of the quality of the reconstruction. As the SBR decreases, we observe the higher uncertainty values. This experiment shows the effectiveness of the proposed method in the real dataset with an accurate reconstruction having uncertainty estimation. The proposed method takes 1.71 seconds in total — 1.53 seconds for initial multiscale estimation and 0.18 seconds for inference. The memory usage is 2.5GB for initial multiscale estimation and 0.5GB for inference. Compared to the Art scene, the memory usage is much lower due to the smaller spatial resolution.
We conduct an additional experiment on another real dataset provided in [25]. The histograms in this dataset have a spatial-time resolution of $256\times 256\times 1536$, with scenes recorded in challenging environments. To make dual peaks in each pixel’s histogram, we modify the dataset by duplicating the portion of the histogram corresponding to the object. Fig. 9 shows the reconstructed pointclouds. In the checkerboard scene (1st row), both BU3D and our method show noisy results but produce flatter surfaces than Peng. In contrast, RT3D yields inaccurate sparse point clouds due to irregular reflectivity. In the elephant scene (2nd row), Peng and our method show fewer artifacts than BU3D. In the last row, the proposed method observes some outliers, but produces less noisy results than BU3D. Considering the fact that the proposed method runs only once, it shows a competitive performance compared to Peng’s results.
@@ VI Conclusion
We have proposed a novel deep learning method for dual-peak reconstruction in single-photon Lidar imaging. Unlike existing deep learning-based methods, the proposed method supports the dual-peak reconstruction in a single run with uncertainty information. The experimental results show that the proposed method presents less artifacts such as bleeding effects with accurate results, although the proposed method often observes outliers in low photon cases. The proposed method can be extended to operate on unstructured or arbitrarily sampled spatial coordinates, such as measurements obtained from arbitrary directions. Adapting the method to such cases would involve modifying the loss computation from grid-based to point-based such as Chamfer distance. A limitation of the current method is to assume a fixed number of peaks per pixel, constrained to a dual-peak model. Future research directions include extending the framework from dual-peak to multi-peak reconstruction scenarios and incorporating reflectivity estimation capabilities. Another future work can be to develop adaptive mechanisms to automatically detect and handle varying numbers of targets per pixel.